7. What voltage will be generated by a cell that consists of an iron electrode immersed in 0.5M FeSO4 solution and a copper electrode immersed in 1M CuSO4solution at 298 K. Given EoFe= -0.41V and EoCu=0.34V. Write the cell representation and cell reaction.
1
Expert's answer
2019-09-12T03:31:30-0400

Cell representation


"Fe|Fe^{2+}(0.5M)||Cu^{2+}(1M)|Cu"

We know that

"E_{cell}=E^0_{cell}-\frac{0.0591}{n}log\frac{[Fe^{2+}]}{[Cu^{}2+]}"

"n=2"

"[Fe^{2+}]=0.5M"

"[Cu^{2+}]=1M\\E^0_{cell}=E^0_{cathode}-E^0_{anode}=0.34-(-.41)\ V=.75\ V"

Hence

"E_{cell}=""0.75+0.008895=0.758895\ V"